A radiator contains 15 quarts of fluid, 30% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 45% antifreeze? (Round your answer to two decimal places.)
Answer by richwmiller(17219)

0.3*15-0.3x+1x=0.45(15)
4.5+0.7x=6.75
0.7x=6.75-4.5
0.7x=2.25
x=3.21428571 rounded 3.21
check
0.3*11.7857143+1*3.21428571=0.45(15)
3.53571429+3.21428571=6.75
6.75=6.75
ok
coderep